# Tidal Stretching and Compression in Black Bounce Backgrounds

### v1

@@ I Introduction
Another way to address the problem of singularities in solutions is through spacetimes known as black bounces (BBs). In this type of spacetime, the singularity and its surroundings are removed and replaced by the throat of a wormhole [54]. The first BB model was proposed by Hochberg and Visser in 1997 and was later further studied by Simpson and Visser in 2018 [55, 56]. This metric, known as the Simpson-Visser (SV) spacetime, can be obtained through a regularization process of the Schwarzschild metric by introducing the parameter $a$, which regularizes the spacetime by replacing $x\rightarrow\sqrt{x^{2}+a^{2}}$ with the metric coefficients. Depending on the value of this parameter, the solution can represent a singular BH, a regular BH, or a wormhole. From the perspective of BH shadows, these solutions can closely mimic their singular counterparts [57, 58, 59, 60]. Several BB models are constructed using this regularization method [61, 62, 63, 64, 65, 66, 67]; however, other models arise from different approaches [68, 69, 70, 71, 72, 73, 74, 75].
@@
In the coordinate system where the radial coordinate can take values between $-\infty<x<+\infty$ (including zero), BB geometries generally have their line element written in the form of
@@ III Geodesics
$-1=-f(r)\dot{t}^{2}+\frac{\dot{r}^{2}}{h(r)}+r^{2}\dot{\theta}^{2}+r^{2}\sin^{2}\theta\dot{\phi}^{2},$ (14)
@@ IV.1 General equations
[…]
@@
[…]
@@
where $\xi^{\hat{\beta}}(b)$ is the $\hat{\beta}=\{\hat{r},\hat{\theta},\hat{\phi}\}$ component of the infinitesimal displacement vector at $r=b$. Furthermore, condition (ICI) corresponds to the release from rest at $r=b$, a body composed of dust without internal interaction, while condition (ICII) corresponds to letting a body composed of dust explode at $r=b$. In this work, we will focus only on the condition (ICI).
@@ IV.2 Tidal Forces effects in SV spacetimes
where $\,{}_{2}F_{1}\left(x_{1},x_{2};x_{3};x_{4}\right)$ is the hypergeometric function and the integration constants $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ can be determined by imposing the boundary conditions at $b$. However, these components of the displacement vector become clearer when analyzed graphically.
@@ IV.3 Tidal Forces effects in Bardeen-type BB spacetimes
where the integration constants $c_{1}$, $c_{2}$, $d_{1}$, and $d_{2}$ can be determined by imposing the boundary conditions in $b$. Although the expressions are not written in terms of special functions, as in the SV case, the analytical expressions are neither clear nor informative. Therefore, it is more appropriate to analyze the components of the displacement vector graphically.
@@ IV.4 Tidal Forces effects in holonomy corrected BH
where the constants $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are determined by imposing the appropriate boundary conditions. Although analytical expressions are simple, the best way to extract information from these solutions is through graphical representations.

### v2

@@ I Introduction
Another way to address the problem of singularities in solutions is through spacetimes known as black bounces (BBs). In this type of spacetime, the singularity and its surroundings are removed and replaced by the throat of a wormhole [54]. The first BB model was proposed by Hochberg and Visser in 1997 and was later further studied by Simpson and Visser in 2018 [55, 56]. This metric, known as the Simpson-Visser (SV) spacetime, can be obtained through a regularization process of the Schwarzschild metric by introducing the parameter $a$, which regularizes the spacetime by replacing $x\rightarrow\sqrt{x^{2}+a^{2}}$ with the metric coefficients. Depending on the value of this parameter, the solution can represent a singular BH, a regular BH, or a wormhole. From the perspective of BH shadows, these solutions can closely mimic their singular counterparts [57, 58, 59, 60]. Several BB models are constructed using this regularization method [61, 62, 63, 64, 65, 66, 67] (for general aspects and consistency, see Ref. [68]); however, other models arise from different approaches [69, 70, 71, 72, 73, 74, 75, 76].
@@
In the coordinate system where the radial coordinate can take values between $\infty<x<+\infty$ (including zero), BB geometries generally have their line element written in the form of
@@ III Geodesics
$1=-f(r)\dot{t}^{2}+\frac{\dot{r}^{2}}{h(r)}+r^{2}\dot{\theta}^{2}+r^{2}\sin^{2}\theta\dot{\phi}^{2},$ (14)
@@ IV.1 General equations
where the hats indices are the tetrad basis indices, and $\eta_{\hat{a}\hat{b}}=\text{diag}(-1,\;1,\;1,\;1)$ is the Minkowski metric in Cartesian coordinates.
@@
$\hat{e}_{\hat{0}}^{\;\;\mu}$ $=$ $\left(\frac{E}{f},\;-\sqrt{\frac{(E^{2}-f)h}{f}},\;0,\;0\right),$ (21)
$\hat{e}_{\hat{1}}^{\;\;\mu}$ $=$ $\left(-\frac{E^{2}-f}{f},\;E\sqrt{\frac{h}{f}},\;0,\;0\right),$ (22)
$\hat{e}_{\hat{2}}^{\;\;\mu}$ $=$ $\left(0,\;0,\;\frac{1}{r},\;0\right),$ (23)
$\hat{e}_{\hat{3}}^{\;\;\mu}$ $=$ $\left(0,\;0,\;0,\;\frac{1}{r\sin\theta}\right).$ (24)
@@
where $\xi^{\hat{\beta}}(b)$ is the $\hat{\beta}=\{\hat{r},\hat{\theta},\hat{\phi}\}$ component of the infinitesimal displacement vector at $r=b$. Furthermore, condition (ICI) corresponds to the release from rest at $r=b$, a body composed of dust without internal interaction, while condition (ICII) corresponds to letting a body composed of dust explode at $r=b$. In this work, although we mention both conditions (ICI) and (ICII) for completeness, we restrict our analysis to (ICI), since it corresponds to the physically relevant situation of a body released from rest at $r=b$, which is the main focus of this study.
@@ IV.2 Tidal Forces effects in SV spacetimes
where $_{2}F_{1}\left(x_{1},x_{2};x_{3};x_{4}\right)$ is the hypergeometric function and the integration constants $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ can be determined by imposing the boundary conditions at $b$. These analytical expressions represent the special case where $b\to\infty$. Next, we shall analyze graphically the behavior of the displacement vector by considering a fixed value of $b=100m$.
@@ IV.3 Tidal Forces effects in Bardeen-type BB spacetimes
where the integration constants $c_{1}$, $c_{2}$, $d_{1}$, and $d_{2}$ can be determined by imposing the boundary conditions in $b$. Although the expressions are not written in terms of special functions, as in the SV case, the analytical expressions are neither clear nor informative. Once again, these analytical expressions represent the special case where $b\to\infty$ and next, we shall analyze graphically the behavior of the displacement vector by considering a fixed value of $b=200m$.
@@ IV.4 Tidal Forces effects in holonomy corrected BH
where the constants $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are determined by imposing the appropriate boundary conditions. Next, we plot the behavior of the displacement vector considering the value $b=200m$.